Nitric acid is formed by the reaction of nitrogen dioxide with water: 3NOβ‚‚(𝑔) + Hβ‚‚O(β„“) β†’ NO(𝑔) + 2HNO₃(π‘Žπ‘ž) How many grams of nitrogen monoxide, NO, are produced when 180.2 g of Hβ‚‚O react with nitrogen dioxide?


1
Expert's answer
2021-06-02T06:31:18-0400

First we must calculate the amount (in moles) of water:


"n(H_{2}O)=\frac{m(H_{2}O)}{M(H_{2}O)}=\frac{180.2g}{18.02g/mol}=10.00mol;"


Now, the amount (in moles) of nitrogen monoxide would be the same as that of water reacted:


"n(NO)=n(H_{2}O)=10.00mol;"


Finally, we can calculate the mass of the nitrogen monoxide formed:


"m(NO) = n(NO)\cdot M(NO)=10.00mol\cdot 30.01g/mol=300.1g."